OpenStudy (anonymous):

Divide: (3x-15)/(14y) ÷ (x^2-15x+50)/7

OpenStudy (anonymous):

\[((3x-15)\div14y)(7/((x-10)(x-5)))\]

OpenStudy (anonymous):

extract the 3 from the 3x-15 which gives u 3(x-5)

OpenStudy (anonymous):

\[\frac{ 3x-15 }{ 14y }\div \frac{ x^2-15x+50 }{ 7 }\]

OpenStudy (anonymous):

now u can simplify more

OpenStudy (anonymous):

\[\frac{ 3(x-5) }{ 14y }\times \frac{ 7 }{ x^2-15x+50 }\] I also went ahead and did the recipricle... which I chould've done to begin with

OpenStudy (anonymous):

yup then factorize the x^2-15x+50

OpenStudy (anonymous):

to (x-5)(x-10)

OpenStudy (anonymous):

\[\frac{ 3(x-5) }{ 14y }\times \frac{ 7 }{ (x-5)(x-10) }\]

OpenStudy (anonymous):

yup now cancel the x-5 and 7/14 is 1/2

OpenStudy (anonymous):

so in the end you should get 3/(2y(x-10))

OpenStudy (anonymous):

yea, that's what I got. Thank you!